Solve the equation. Check for extraneous solutions.
4|5-6x|=3x+4

Answer by lwsshak3(11628)

Solve the equation. Check for extraneous solutions.
4|5-6x|=3x+4
**
solve for two conditions:
4|5-6x|=3x+4
4(5-6x)=3x+4
20-24x=3x+4
27x=16
x=16/27
..
4*-(5-6x)=3x+4
-20+24x=3x+4
21x=24
x=24/21
solution: x=16/27 or 24/21
..
Check:
For x=16/27
4|5-6x|=3x+4
4|5-6*16/27|=3*16/27+4
5.778=5.778
..
for x=24/21
4|5-6x|=3x+4
4|5-6*24/21|=3*24/21+4
7.4286=7.4286
..
Check ok, no extraneous solutions.
solution: x=16/27 or 24/21